OpenStudy (anonymous):

You deposit $7,900 in a money-market account that pays an annual interest rate of 4.3%. The interest is compounded quarterly. How much money will you have after 3 years?

OpenStudy (anonymous):

have you seen the 'single payment' compounding equation that looks like this: \[FV = PV(1+i)^n\] 'FV = PV(1+i)^n

OpenStudy (anonymous):

no I was only given A=P(1+r/k)^n*k

OpenStudy (anonymous):

ok, we will use that one then

OpenStudy (anonymous):

I have nooo idea how to use it

OpenStudy (anonymous):

we are solving for A, A is the future amount A=? P is the present value, so we were told 7,900, so p=7900 r is the interest rate, so 4.3%, but in decimal form so r=0.043 k is the how often it is 'compounded', it says quarterly, so k=4 n is how long, 3 years, so n=3 \[A=P(1+ \frac {r}{k})^{n*k}\] plug in the values

OpenStudy (anonymous):

A=7900(1+0.043/4)^3*4 would you use the order of operations?

OpenStudy (anonymous):

good work ^_^ yes, use order of operations, A=7900 ( 1+ (0.043/4) )^ (3*4)

OpenStudy (anonymous):

A=7900 (1+ (0.01075) )^ (12) I feel like I did something wrong..

OpenStudy (anonymous):

you are good, nothing worng, keep finishing out the math ^_^

OpenStudy (anonymous):

A=7900 (1.01075)^12 now do I go from left to right or..?

OpenStudy (anonymous):

when you get to here A=7900 (1.01075)^12 you do this part first: A=7900 ( (1.01075)^12 ) always do the power before the timesing

OpenStudy (anonymous):

okay, so A=7900 (12.129) A=95,819.1 right?

OpenStudy (anonymous):

close, i think there was some mix-up with the power. we had this: A=7900 ( (1.01075)^12 ) becomes A=7900(1.1369)

OpenStudy (anonymous):

well dang. so, A=8981.51

OpenStudy (anonymous):

correct ^_^

OpenStudy (anonymous):

:D thank you!

OpenStudy (anonymous):

yep ^_^